A dodecagon is a polygon with 12 sides. How many degrees is each of the interior angles in a regular dodecagon? Show or explain how you got your answer.
Answer by jim_thompson5910(23765)

I = (180(n-2))/n

I = (180(12-2))/(12)

I = (180(10))/(12)

I = 1800/(12)

I = 150


So the measure of each interior angle of a 12-sided polygon is 150 degrees.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim